OpenStudy (mtalhahassan2):

A child throws a ball with a velocity of 1.93 m/s [horizontally] at a height of 0.75 m. If air resistance is negligible, the horizontal range of the ball is: Question 6 options: A) 0.30 m B) 4.9 m C) 0.76 m D) 0.20 m E) 0.73 m

OpenStudy (mtalhahassan2):

@Astrophysics @Vincent-Lyon.Fr

OpenStudy (mtalhahassan2):

@study_buddy99 @yaya090600 @ltrout @Sienna_Kate

OpenStudy (mtalhahassan2):

@ganeshie8 @Preetha @ParthKohli @Kainui @Kkutie7

OpenStudy (mtalhahassan2):

@baru @Irrati0nal

OpenStudy (baru):

first find time(t) required for an object to free fall from a height of 0.75m (use s=1/2gt^2)

OpenStudy (baru):

1.93*t gives the required answer

OpenStudy (mtalhahassan2):

g be 9.8 m/s right?

OpenStudy (baru):

yes

OpenStudy (mtalhahassan2):

well i am getting 18.25

OpenStudy (baru):

\[0.75=0.5 \times 9.8 \times t^2\]

OpenStudy (baru):

t=0.39

OpenStudy (mtalhahassan2):

so for finding the time you divide 0.75/1.93 which gives 0.39 right?

OpenStudy (baru):

no

OpenStudy (baru):

we get 't' from solving this equation\[0.75=0.5 \times 9.8 \times t^2\]

OpenStudy (baru):

\[t=\sqrt{\frac{0.75}{0.5\times 9.8} }\]

OpenStudy (mtalhahassan2):

why did you put square root?

OpenStudy (baru):

after finding 't' multiply 1.93 X t to find "range"

OpenStudy (baru):

the formula has t square, (\(t^2\))

OpenStudy (mtalhahassan2):

oh yeah

OpenStudy (mtalhahassan2):

so know we have to multiply 1.93x0.39

OpenStudy (baru):

yes

OpenStudy (mtalhahassan2):

i get 0.7527

OpenStudy (baru):

option C is the closest match

OpenStudy (mtalhahassan2):

yeah

OpenStudy (mtalhahassan2):

@baru thanks a lot

OpenStudy (baru):

no problem :)